# Physics-informed machine learning for combustion: A review

### v2

@@ Abstract
Physics-informed machine learning (PIML) represents an emerging paradigm that integrates various forms of physical knowledge into machine learning (ML) components, thereby enhancing the physical consistency of ML models compared to purely data-driven paradigms. The field of combustion, characterized by a rich foundation of physical laws and abundant data, is undergoing a transformation due to PIML. This paper aims to provide a comprehensive overview of PIML for combustion, systematically outlining fundamental principles, significant contributions, key advancements, and available resources. The application of PIML in combustion is categorized into three domains: combustion chemical kinetics, combustion reacting flows, and other combustion-related problems. Additionally, current challenges, potential solutions, and practical guidelines for researchers and engineers will be discussed. A primary focus of this review is to demonstrate how combustion laws can be integrated into ML, either through soft or hard constraints, via loss functions or representation models, and within coordinate-to-variable or field-to-field paradigms. This paper shows that PIML offers a unified framework linking physics, model, and data in combustion–integrating physical knowledge in model-to-data simulation and reconstruction tasks, as well as data-to-model modeling tasks–resulting in enhanced data, improved physical models, and more reliable ML models. PIML for combustion presents significant opportunities for both the combustion and ML communities, encouraging greater collaboration and cross-disciplinary engagement.
Keywords physics-informed machine learning (PIML)  $\cdot$ combustion  $\cdot$ chemical kinetics  $\cdot$ reacting flow
@@ 2.2 Reviews and scoping
Some reviews focus on the PIML for fluid mechanics 90, 91, 92, which can provide some insights for PIML for combustion flows. Note that the classification of PIML in Reference 91 is again consistent with the References 68, 83, 85, 89, including the physics-informed features and labels, architectures, and loss functions. However, these reviews rarely include combustion or reacting flows. Recently, Sophiya et al. 93 reviewed the PINNs in industrial gas turbines (IGTs) research, highlighting their diverse applications in aerodynamics and aeromechanics. However, there are few combustion applications introduced. A more combustion-related review is Reference 94, which focuses on the PIML in multiphysics modeling in chemical engineering. The applications of PIML in momentum transfer/fluid flow, heat transfer, mass transfer, and chemical reaction were first reviewed separately, then the PIML for multiphysics coupling, surrogate model construction, and design optimization were presented. PIML for chemical engineering has also been briefly reviewed in Reference 87. However, it should be noted that combustion and chemical engineering do not contain each other, but intersect with each other. They both involve chemical reactions but the range of combustion is narrower: a high-temperature exothermic redox chemical reaction by definition. On the other hand, combustion is not limited to chemical engineering, but also involves power generation, transportation, fire safety, and others. Deng et al. 47 reviewed the SciML in combustion, presenting how SciML can assist: (1) modeling, such as in parameter estimation and reaction mechanism generation, (2) simulation, such as via DL and knowledge-guided ML, and (3) reconstruction, such as for sparse reconstruction and mutual inference. However, this review does not focus on PIML and does not elaborate on how physics and ML can be integrated in diverse ways, which are the main points of this paper.
@@ 4.1 Solving ODEs of chemical kinetics
Ji et al. 261 first proposed the stiff-PINNs (shown in Figure 5), using PINNs to solve stiff chemical kinetic ODEs, where the quasi-steady-state assumption (QSSA) was employed to reduce the stiffness of the system. In QSSA, QSS species are identified by assuming that their net production rates are negligible compared to their consumption and production rates, then the ODEs of QSS species become differential-algebraic equations (DAEs). For example, $B$ can be chosen as the QSS species in the ROBER problem, so its algebraic equation can be obtained by assuming $\mathrm{d}y_{2}/\mathrm{d}t=0$. Because QSS species are usually radicals and unstable intermediates with short time scales, the elimination of their ODEs can greatly reduce the stiffness. The stiff-PINNs were used to solve the ROBER and POLLU problems with one and ten QSS species, respectively, and yielded satisfactory results. De Florio et al. 262 used X-TFC 136 (see Section 2.3.2) to solve stiff chemical kinetic ODEs, where the TFC can guarantee the ICs are exactly satisfied, as illustrated in Figure 6. The temporal domain was decomposed into several intervals to deal with the long-time issue. The X-TFC showed great performance for 4 tested problems, and even outperformed the traditional numerical methods in some cases. Wang et al. 263 used PINNs to solve the ROBER problem. Hard constraints on ICs via logarithmic functions were employed, yielding accurate results. Galaris et al. 264 proposed the physics-informed random projection NNs (PI-RPNNs) to solve stiff ODEs, where the single-layer NN had fixed weights between the input and hidden layer so only the parameters from the hidden to the output layer needed to be optimized. The loss functions were minimized by the Gauss-Newton method and the hard constraints on ICs were designed. Four stiff ODE problems were tested and PI-RPNN yielded good accuracy and can outperform the traditional solvers in the cases with steep gradients. Later, Fabiani et al. 265 extended the PI-RPNNs to other stiff ODEs and three index-1 DAEs. For more practical applications, Li et al. 266 applied PINNs to solve the coal gasification chemical kinetic problems. The unreacted-core shrinking model was adopted for the gas-solid reactions, the temperature was simplified to be constant, and the ICs were hardcoded. Results showed that PINNs could accurately predict the production rates of the gaseous species.
Figure 5: Schematic of the stiff-PINN for the ROBER problem. Reprinted from Reference 261, with permission.
@@
Figure 6: Schematic of the X-TFC algorithm to solve stiff chemical kinetics equations. Reprinted from Reference 262, with permission.
@@
Although the above studied focus on solving the chemical kinetic ODEs in a forward way without data supervision, inverse problems can also be solved when the available data are utilized. Weng et al. 267 developed the multiscale PINNs for stiff chemical kinetic problems, where the species with different time scales were classified and trained by multiple NNs. By introducing ground truth data points, the proposed method showed superior ability on solving three stiff chemical kinetic problems. Gusmao et al. 268 proposed the kinetics-informed NNs (KINNs) to solve forward and inverse problems of chemical kinetics, as shown in Figure 7. Practical species and reaction types were considered, including three types of species (gas, absorbed molecules, and surface radicals) and five types of reactions (denoted by g, a, d, c, and s). ICs were enforced by an IC operator coupled with a shallow NN and species conservation was enforced by trigonometric transformations. In forward problems, no data were provided, while in inverse problems, synthetic data were available and three scenarios were designed based on the information about chemical intermediates and adsorbed molecules, and the presence of noise. The unknown kinetic parameters were optimized together with the NN parameters in inverse problems. The results showed that KINNs can accurately solve the ODEs in forward problems and estimate the kinetic parameters in inverse problems. Later, KINNs were applied to model the data from temporal analysis of products 286 and extended to a robust version by incorporating the maximum likelihood approach 287. Bibeau et al. 269 used PINNs to predict the kinetics of biodiesel production in microwave reactors. The system contained five species and three reactions. The Arrhenius equation was linearized and the temperature evolution was simplified to a first-order ODE, resulting in 15 unknown parameters related to the reaction kinetics and energy balance, which were estimated by PINNs from experimental data. Results showed that the reaction kinetics of the biodiesel process could be identified successfully by PINNs with satisfactory extrapolation ability.
Figure 7: Kinetics-informed neural networks. Reprinted from Reference 268, with permission.
@@ 4.2 Surrogate models of chemical kinetics
Almeldein et al. 270 developed the physics-informed surrogate NNs for accelerating chemical kinetics calculations following Equation (4.12), where the conservation of four elements was added in the loss function. The mixture-of-experts architecture was adopted to allow multiple subnetworks to specialize in different regimes, such as pre-ignition low-temperature chemistry, high-temperature reactions, and post-combustion equilibriums. The 0D CH4-air combustion was tested, demonstrating the effectiveness of element conservation and the acceleration ability of the surrogate models on both CPUs and GPUs. Zhang et al. 271, 288 proposed the parametric physics-informed surrogate NNs for combustion reaction kinetics (CRK-PINNs) based on the paradigm of Equation (4.12). As illustrated in Figure 8, the kinetic ODEs and the conservations of elements, species, and energy were considered in the loss function to improve the physical completeness and low data dependence. A logarithmic normalization strategy was adopted to mitigate the multiscale of species and stiffness of the system. Based on the H2-air mechanism with 10 species and 21 reactions, the CRK-PINN was validated in three scenarios: solving forward and inverse problems, surrogate chemical kinetics, and accelerating simulations of combustion reacting flows. In the first scenario, CRK-PINN can solve the ignition process without data, reconstruct the process with sparse data, and infer intermediate species and temperature with the data of major species. In the second scenario, CRK-PINN can surrogate chemical kinetics in wide operating conditions and exhibited higher accuracy and lower data dependence than data-driven NNs for both time-continuous and time-advanced predictions. In the last scenario, CRK-PINN can replace the direct integration solver to predict the reaction source terms in the DNS of a 2D laminar Bunsen flame and a 3D turbulent jet flame, showing superior accuracy than data-driven NNs and great acceleration compared to the direct integration method.
Figure 8: CRK-PINN architecture. Reprinted from Reference 271, with permission.
@@
which is a good alternative to the logarithmic normalization 271. Although the paradigm of Equation (4.13) was adopted, only the species were directly predicted by the NNs, while the temperature and pressure were updated by transport fluxes, thus ensuring the mass and energy conservation. The algorithm has been validated on various CFD cases of combustion and been applied to DeepFlame 118, a deep learning empowered open-source platform for reacting flow simulations. Wang et al. 116 extended the method by imposing constraints on element conservations in both soft and hard ways. As illustrated in Figure 9, the ANN-soft penalized the element conservations by adding a loss term, while ANN-hard enforced the element conservations by adding a correction term, $\Delta\mathbf{Y}_{c}$, to the NN-predicted mass fractions. The correction term can be easily derived by the element conservations and the changes of species mass fractions, and was first computed in the Box-Cox transformed space to avoid the multiscale issue. One advantage of ANN-hard over ANN-soft is that there is no cumbersome adjustment of the loss weights. In terms of prediction accuracy and error of element conservations, ANN-hard exhibited superior performance than ANN-soft and they all outperformed the regular ANN on various flame cases. The computational overhead caused by the correction term was negligible, demonstrating the efficiency of ANN-hard. In fact, the correction method to ensure element conservations has been adopted in earlier studies 54, 55, where the NNs were trained to surrogate the kinetics of reduced chemistry, following the paradigm of Equation (4.15).s
Figure 9: ANN-hard: schematic and comparisons with other approaches. Reprinted from Reference 116, with permission.
@@
In addition to the paradigm of function learning, some studies follow the paradigm of operator learning to learn combustion chemical kinetics, using neural operators such as the DeepONet 98 and FNO 99. Some data-driven DeepONets have been used to surrogate chemical kinetics 289, 142, 143, where the input of branch nets and the final output are similar to Equation (4.12), while the input of truck nets is related to the temporal coordinates or increments, enabling the models to learn kinetics with variable $\Delta t$, which is often fixed in classical surrogate models. To some extent, they can be viewed as a hybrid paradigm of AR and NAR learning. One insight that these works can provide is that learning kinetics in the latent space with reduced dimension can improve computational efficiency, which can be achieved by integrating a self-supervised autoencoder (AE) 142, 143, as shown in Figure 10. Kumar et al. 143 also accelerated the computation by selecting representative species for kinetics learning. The other scalars can be reconstructed by a pre-trained correlation NN. This strategy was also adopted in Reference 105, 106, where the species and element conservations were integrated into the loss functions of both correlation NN and DeepONet, thus constituting a physics-informed learning. The framework was tested for CH4 oxidation, demonstrating that the physical constraints can reduce the conservation errors without loss of prediction accuracy. Zanardi et al. 107 used the PI-DeepONets to learn the solution operator of non-equilibrium chemistry in hypersonic flows, where the coarse-grained (CG) kinetic ODEs were considered. The input of the branch and truck nets were the ICs and temporal coordinates, respectively. The framework was tested on the O2–O mixture, exhibiting two orders of magnitude in speedup compared to numerical integrators. Later, Zanardi et al. 108 extended the framework to be more physics-informed by many strategies, including a Boltzmann transformation layer to enforce local equilibrium distributions between states in the same group, the hierarchical architecture and transfer learning between different temporal scales, and the adaptive online pruning technique. Weng et al. 272 applied FNO to learn stiff chemical kinetics, following the paradigm of Equation (4.12). The loss of element and species conservations were considered and the BCT (Equation (4.16)) was also adopted. Based on the focal loss 290, a balanced loss function was proposed to address the unbalanced sampling issue in complex reacting flow problems. The results on four test cases (ROBER, POLLU, H2 autoignition, and H2–NH3 turbulent jet flame) demonstrated the effectiveness of physical constraints, BCT, and the balanced loss function.
Figure 10: AE integrated DeepONet for stiff chemical kinetics. Reprinted from Reference 142, with permission.
@@
Some studies used NNs to learn chemical tabulations, which can be viewed as surrogate modeling of chemical kinetics as well. In References 273, 274, 275, NNs were used to learn the tabulation of thermochemistry following the paradigm of Equation (4.13). The hybrid flamelet/random data were generated for training and multiple NNs were used to predict each species with different magnitudes of concentration changes. The species conservation (Equation (3.1)) was enforced by normalizing the predicted species concentrations. Readshaw et al. 276 extended this framework by imposing constraints on element conservation, which was achieved by recalculating the mass fractions of selected residual species after the initial prediction. The results demonstrated that this conserved NN exhibited higher accuracy and physical consistency than the non-conserved NNs. Salunkhe et al. 277 developed the framework of ChemTab to jointly learn the progress variable generation (PVG) and manifold approximation (MA). For the PVG, the high dimensional species mass fractions are linearly embedded to the progress variables: $\mathbf{C_{pv}}=\mathbf{YW}$, while MA means the mapping between $\mathbf{C_{pv}}\oplus Z$ and the source terms is learned by a NN, where $Z$ is the mixture fraction. Therefore, the overall framework is similar to Equation (4.15). Inspired by the principal component analysis (PCA), some constraints were added in the loss function, including the orthogonality of $\mathbf{W}$ and $\mathbf{C_{pv}}\oplus Z$, and the unit-norm constraint on $\mathbf{W}$. It was demonstrated that removing any of these constraints will make the results worse. Salunkhe et al. 278 extended the ChemTab by adding another NN to learn the relationship between $\mathbf{C_{pv}}\oplus Z$ and lower dimensional source terms, which was computed using the same projection method and the same $\mathbf{W}$, as illustrated in Figure 11. The unit-norm constraint on $\mathbf{W}$ was replaced with the non-negativity constraint, which was achieved in a hard manner by truncating the negative values.
Figure 11: ChemTab extended architecture and training procedure. Reprinted from Reference 278, with permission.

### v3

@@ Abstract
Physics-informed machine learning (PIML) represents an emerging paradigm that integrates various forms of physical knowledge into machine learning (ML) components, thereby enhancing the physical consistency of ML models compared to purely data-driven paradigms. The field of combustion, characterized by a rich foundation of physical laws and abundant data, is undergoing a transformation due to PIML. This paper aims to provide a comprehensive overview of PIML for combustion, systematically outlining fundamental principles, significant contributions, key advancements, and available resources. The application of PIML in combustion is categorized into three domains: combustion chemical kinetics, combustion reacting flows, and other combustion-related scenarios. Additionally, current challenges, potential solutions, and practical guidelines for researchers and engineers will be discussed. A primary focus of this review is to demonstrate how combustion laws can be integrated into ML, either through soft or hard constraints, via loss functions or representation models, and within coordinate-to-variable or field-to-field paradigms. This paper shows that PIML offers a unified framework linking physics, model, and data in combustion–integrating physical knowledge in model-to-data simulation and reconstruction tasks, as well as data-to-model modeling tasks–resulting in enhanced data, improved physical models, and more reliable ML models. PIML for combustion presents significant opportunities for both the combustion and ML communities, encouraging greater collaboration and cross-disciplinary engagement.
Keywords physics-informed machine learning $\cdot$ physics-informed neural network $\cdot$ combustion $\cdot$ chemical kinetics $\cdot$ reacting flow
@@ 2.2 Reviews and scoping
Some reviews focus on the PIML for fluid mechanics 90, 91, 92, which can provide some insights for PIML for combustion flows. Note that the classification of PIML in Reference 91 is again consistent with the References 68, 83, 85, 89, including the physics-informed features and labels, architectures, and loss functions. However, these reviews rarely include combustion or reacting flows. Recently, Sophiya et al. 93 reviewed the PINNs in industrial gas turbines (IGTs) research, highlighting their diverse applications in aerodynamics and aeromechanics. However, there are few combustion applications introduced. A more combustion-related review is Reference 94, which focuses on the PIML in multiphysics modeling in chemical engineering. The applications of PIML in momentum transfer/fluid flow, heat transfer, mass transfer, and chemical reaction were first reviewed separately, then the PIML for multiphysics coupling, surrogate model construction, and design optimization were presented. PIML for chemical engineering has also been briefly reviewed in Reference 87. However, it should be noted that combustion and chemical engineering do not contain each other, but intersect with each other. They both involve chemical reactions but the range of combustion is narrower: a high-temperature exothermic redox chemical reaction by definition. On the other hand, combustion is not limited to chemical engineering, but also involves power generation, transportation, fire safety, and others. Deng et al. 47 reviewed the SciML in combustion, presenting how SciML can assist: (1) modeling, such as in parameter estimation and reaction mechanism generation, (2) simulation, such as via DL and knowledge-guided ML, and (3) reconstruction, such as for sparse reconstruction and mutual inference. Compared to this review, our review focuses more on PIML, elaborating on how physics and ML can be integrated in diverse ways.
@@ 4.1 Solving ODEs of chemical kinetics
Figure 5 shows some studies on PIML for solving ODEs of chemical kinetics. Ji et al. 261 first proposed the stiff-PINNs, using PINNs to solve stiff chemical kinetic ODEs, where the quasi-steady-state assumption (QSSA) was employed to reduce the stiffness of the system. In QSSA, QSS species are identified by assuming that their net production rates are negligible compared to their consumption and production rates, then the ODEs of QSS species become differential-algebraic equations (DAEs). For example, $B$ can be chosen as the QSS species in the ROBER problem, so its algebraic equation can be obtained by assuming $\mathrm{d}y_{2}/\mathrm{d}t=0$. Because QSS species are usually radicals and unstable intermediates with short time scales, the elimination of their ODEs can greatly reduce the stiffness. The stiff-PINNs were used to solve the ROBER and POLLU problems with one and ten QSS species, respectively, and yielded satisfactory results. De Florio et al. 262 used X-TFC 136 (see Section 2.3.2) to solve stiff chemical kinetic ODEs, where the TFC can guarantee the ICs are exactly satisfied, as illustrated in Figure 5(b). The temporal domain was decomposed into several intervals to deal with the long-time issue. The X-TFC showed great performance for 4 tested problems, and even outperformed the traditional numerical methods in some cases. Wang et al. 263 used PINNs to solve the ROBER problem. Hard constraints on ICs via logarithmic functions were employed, yielding accurate results. Galaris et al. 264 proposed the physics-informed random projection NNs (PI-RPNNs) to solve stiff ODEs, where the single-layer NN had fixed weights between the input and hidden layer so only the parameters from the hidden to the output layer needed to be optimized. The loss functions were minimized by the Gauss-Newton method and the hard constraints on ICs were designed. Four stiff ODE problems were tested and PI-RPNN yielded good accuracy and can outperform the traditional solvers in the cases with steep gradients. Later, Fabiani et al. 265 extended the PI-RPNNs to other stiff ODEs and three index-1 DAEs. For more practical applications, Li et al. 266 applied PINNs to solve the coal gasification chemical kinetic problems. The unreacted-core shrinking model was adopted for the gas-solid reactions, the temperature was simplified to be constant, and the ICs were hardcoded. Results showed that PINNs could accurately predict the production rates of the gaseous species.
@@
Figure 5: PIML for solving ODEs of chemical kinetics. (a) Stiff-PINN for the ROBER problem 261. (b) Using the X-TFC algorithm to solve stiff chemical kinetics equations 262. (c) Kinetics-informed neural networks 268. Figures are reprinted with permission.
@@
Although the above studied focus on solving the chemical kinetic ODEs in a forward way without data supervision, inverse problems can also be solved when the available data are utilized. Weng et al. 267 developed the multiscale PINNs for stiff chemical kinetic problems, where the species with different time scales were classified and trained by multiple NNs. By introducing ground truth data points, the proposed method showed superior ability on solving three stiff chemical kinetic problems. Gusmao et al. 268 proposed the kinetics-informed NNs (KINNs) to solve forward and inverse problems of chemical kinetics, as shown in Figure 5(c). Practical species and reaction types were considered, including three types of species (gas, absorbed molecules, and surface radicals) and five types of reactions (denoted by g, a, d, c, and s). ICs were enforced by an IC operator coupled with a shallow NN and species conservation was enforced by trigonometric transformations. In forward problems, no data were provided, while in inverse problems, synthetic data were available and three scenarios were designed based on the information about chemical intermediates and adsorbed molecules, and the presence of noise. The unknown kinetic parameters were optimized together with the NN parameters in inverse problems. The results showed that KINNs can accurately solve the ODEs in forward problems and estimate the kinetic parameters in inverse problems. Later, KINNs were applied to model the data from temporal analysis of products 286 and extended to a robust version by incorporating the maximum likelihood approach 287. Bibeau et al. 269 used PINNs to predict the kinetics of biodiesel production in microwave reactors. The system contained five species and three reactions. The Arrhenius equation was linearized and the temperature evolution was simplified to a first-order ODE, resulting in 15 unknown parameters related to the reaction kinetics and energy balance, which were estimated by PINNs from experimental data. Results showed that the reaction kinetics of the biodiesel process could be identified successfully by PINNs with satisfactory extrapolation ability.
@@ 4.2 Surrogate models of chemical kinetics
Some studies on PIML for surrogate modeling of chemical kinetics are illustrated in Figure 6. Almeldein et al. 270 developed the physics-informed surrogate NNs for accelerating chemical kinetics calculations following Equation (4.12), where the conservation of four elements was added in the loss function. The mixture-of-experts architecture was adopted to allow multiple subnetworks to specialize in different regimes, such as pre-ignition low-temperature chemistry, high-temperature reactions, and post-combustion equilibriums. The 0D CH4-air combustion was tested, demonstrating the effectiveness of element conservation and the acceleration ability of the surrogate models on both CPUs and GPUs. Zhang et al. 271, 288 proposed the parametric physics-informed surrogate NNs for combustion reaction kinetics (CRK-PINNs) based on the paradigm of Equation (4.12). As illustrated in Figure 6(a), the kinetic ODEs and the conservations of elements, species, and energy were considered in the loss function to improve the physical completeness and low data dependence. A logarithmic normalization strategy was adopted to mitigate the multiscale of species and stiffness of the system. Based on the H2-air mechanism with 10 species and 21 reactions, the CRK-PINN was validated in three scenarios: solving forward and inverse problems, surrogate chemical kinetics, and accelerating simulations of combustion reacting flows. In the first scenario, CRK-PINN can solve the ignition process without data, reconstruct the process with sparse data, and infer intermediate species and temperature with the data of major species. In the second scenario, CRK-PINN can surrogate chemical kinetics in wide operating conditions and exhibited higher accuracy and lower data dependence than data-driven NNs for both time-continuous and time-advanced predictions. In the last scenario, CRK-PINN can replace the direct integration solver to predict the reaction source terms in the DNS of a 2D laminar Bunsen flame and a 3D turbulent jet flame, showing superior accuracy than data-driven NNs and great acceleration compared to the direct integration method.
Figure 6: PIML for surrogate modeling of chemical kinetics. (a) The architecture of CRK-PINN 271. (b) ANN-hard: schematic and comparisons with other approaches 116. (c) ChemTab extended architecture and training procedure 278. Figures are reprinted with permission.
@@
which is a good alternative to the logarithmic normalization 271. Although the paradigm of Equation (4.13) was adopted, only the species were directly predicted by the NNs, while the temperature and pressure were updated by transport fluxes, thus ensuring the mass and energy conservation. The algorithm has been validated on various CFD cases of combustion and been applied to DeepFlame 118, a deep learning empowered open-source platform for reacting flow simulations. Wang et al. 116 extended the method by imposing constraints on element conservations in both soft and hard ways. As illustrated in Figure 6(b), the ANN-soft penalized the element conservations by adding a loss term, while ANN-hard enforced the element conservations by adding a correction term, $\Delta\mathbf{Y}_{c}$, to the NN-predicted mass fractions. The correction term can be easily derived by the element conservations and the changes of species mass fractions, and was first computed in the Box-Cox transformed space to avoid the multiscale issue. One advantage of ANN-hard over ANN-soft is that there is no cumbersome adjustment of the loss weights. In terms of prediction accuracy and error of element conservations, ANN-hard exhibited superior performance than ANN-soft and they all outperformed the regular ANN on various flame cases. The computational overhead caused by the correction term was negligible, demonstrating the efficiency of ANN-hard. In fact, the correction method to ensure element conservations has been adopted in earlier studies 54, 55, where the NNs were trained to surrogate the kinetics of reduced chemistry, following the paradigm of Equation (4.15).
@@
In addition to the paradigm of function learning, some studies follow the paradigm of operator learning to learn combustion chemical kinetics, using neural operators such as the DeepONet 98 and FNO 99. Some data-driven DeepONets have been used to surrogate chemical kinetics 289, 142, 143, where the input of branch nets and the final output are similar to Equation (4.12), while the input of truck nets is related to the temporal coordinates or increments, enabling the models to learn kinetics with variable $\Delta t$, which is often fixed in classical surrogate models. To some extent, they can be viewed as a hybrid paradigm of AR and NAR learning. One insight that these works can provide is that learning kinetics in the latent space with reduced dimension can improve computational efficiency, which can be achieved by integrating a self-supervised autoencoder (AE) 142, 143. Kumar et al. 143 also accelerated the computation by selecting representative species for kinetics learning. The other scalars can be reconstructed by a pre-trained correlation NN. This strategy was also adopted in Reference 105, 106, where the species and element conservations were integrated into the loss functions of both correlation NN and DeepONet, thus constituting a physics-informed learning. The framework was tested for CH4 oxidation, demonstrating that the physical constraints can reduce the conservation errors without loss of prediction accuracy. Zanardi et al. 107 used the PI-DeepONets to learn the solution operator of non-equilibrium chemistry in hypersonic flows, where the coarse-grained (CG) kinetic ODEs were considered. The input of the branch and truck nets were the ICs and temporal coordinates, respectively. The framework was tested on the O2–O mixture, exhibiting two orders of magnitude in speedup compared to numerical integrators. Later, Zanardi et al. 108 extended the framework to be more physics-informed by many strategies, including a Boltzmann transformation layer to enforce local equilibrium distributions between states in the same group, the hierarchical architecture and transfer learning between different temporal scales, and the adaptive online pruning technique. Weng et al. 272 applied FNO to learn stiff chemical kinetics, following the paradigm of Equation (4.12). The loss of element and species conservations were considered and the BCT (Equation (4.16)) was also adopted. Based on the focal loss 290, a balanced loss function was proposed to address the unbalanced sampling issue in complex reacting flow problems. The results on four test cases (ROBER, POLLU, H2 autoignition, and H2–NH3 turbulent jet flame) demonstrated the effectiveness of physical constraints, BCT, and the balanced loss function.
@@
Some studies used NNs to learn chemical tabulations, which can be viewed as surrogate modeling of chemical kinetics as well. In References 273, 274, 275, NNs were used to learn the tabulation of thermochemistry following the paradigm of Equation (4.13). The hybrid flamelet/random data were generated for training and multiple NNs were used to predict each species with different magnitudes of concentration changes. The species conservation (Equation (3.1)) was enforced by normalizing the predicted species concentrations. Readshaw et al. 276 extended this framework by imposing constraints on element conservation, which was achieved by recalculating the mass fractions of selected residual species after the initial prediction. The results demonstrated that this conserved NN exhibited higher accuracy and physical consistency than the non-conserved NNs. Salunkhe et al. 277 developed the framework of ChemTab to jointly learn the progress variable generation (PVG) and manifold approximation (MA). For the PVG, the high dimensional species mass fractions are linearly embedded to the progress variables: $\mathbf{C_{pv}}=\mathbf{YW}$, while MA means the mapping between $\mathbf{C_{pv}}\oplus Z$ and the source terms is learned by a NN, where $Z$ is the mixture fraction. Therefore, the overall framework is similar to Equation (4.15). Inspired by the principal component analysis (PCA), some constraints were added in the loss function, including the orthogonality of $\mathbf{W}$ and $\mathbf{C_{pv}}\oplus Z$, and the unit-norm constraint on $\mathbf{W}$. It was demonstrated that removing any of these constraints will make the results worse. Salunkhe et al. 278 extended the ChemTab by adding another NN to learn the relationship between $\mathbf{C_{pv}}\oplus Z$ and lower dimensional source terms, which was computed using the same projection method and the same $\mathbf{W}$, as illustrated in Figure 6(c). The unit-norm constraint on $\mathbf{W}$ was replaced with the non-negativity constraint, which was achieved in a hard manner by truncating the negative values.